what is the answer to 7-1/6x=10/3x?
Answer by solver91311(24713)



It depends. Since you didn't use any parentheses, it is impossible to tell whether you meant:



for which you should have written 7 - (1/6)x = (10/3)x

or



for which you should have written 7 - 1/(6x) = 10/(3x)

The two equations have very different solution sets.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism